# perceptual musical features for interpretable audio tagging

### 04 Jan 2024 v2

@@ 1 Introduction
In these cases, the reliability of the training data dictates the reliability of the model, and without interpretability, it is difficult to detect flaws or biases in either data or models [12]. For music especially, where real-world data tend to be diverse, contrary to available datasets [13], it is essential to be able to detect and fix biases to make robust systems that can be applied in real-world settings [14]. As an extreme example, consider a mood tagging system that was trained only on pop and rock music. We cannot expect such a system to be generalized for all music genres, and it would not be possible to predict how it would behave if it were fed for instance classical music. Furthermore, we cannot expect an end-user to know the distribution of the training dataset, and the nuances of training machine learning models, thus such a system would not be applicable in practice, regardless of how well it performed on a test set. Contrarily, if the mood tagging system itself were interpretable, it would be easier for a user to understand why the system assigns specific tags to samples, making it more trustworthy and more likely to be applied in a real-world setting.
@@
In recent years, efforts have emerged aimed at addressing the interpretability challenge within music audio tagging systems [15]. Perceptual musical features are used that are readily comprehensible to human users [16, 17, 18, 19, 20]. These features encompass characteristics imbued with both musical and acoustic significance, such as tonality, rhythmic stability, or loudness, and they exhibit a strong correlation with emotions [21, 22]. Due to this inherent correlation, these features have demonstrated their efficacy in Music Emotion Recognition (MER) tasks. These characteristics are rooted in musical concepts and theoretical knowledge, as noted in a survey [23], though quantitative studies on them remain relatively scarce. In our research, we build upon these concepts and execute a structured approach for achieving interpretable music audio tagging (depicted in Fig. 1), harnessing the power of perceptual features. Our specific contributions can be succinctly summarized as follows:
  • We incorporate three methods for extracting perceptual musical features including signal processing, deep neural networks, and symbolic knowledge.
  • We introduce a novel set of features related to music harmony by setting up a music knowledge base and using reasoning.
  • We interpret model predictions, using four different methodologies, namely SHAP [24], XGBoost [25] feature importance, permutation-based feature importance, and an ablation study.
@@ 2.1 Datasets
The dataset employed in our experiments was the MTG-Jamendo dataset [26], which serves as an open resource for music audio tagging. This dataset was constructed using music accessible through the Jamendo platform, offered under Creative Commons licenses, and enriched with tags provided by content uploaders. Boasting a substantial scale, it encompasses more than 18,486 complete audio tracks, annotated with 56 tags that encompass various mood and theme categories. We extracted MFCCs with 40 coefficients, utilizing a window and FFT size of 2048. Our choice of this specific dataset was predicated on the anticipation that interpretability would offer heightened value in a task involving partially subjective aspects, such as mood detection [27].
We utilized the GTZAN dataset [28] as our secondary dataset. This collection comprises 1000 audio tracks of 30 seconds and features 10 genres, with 100 tracks representing each. The audio files are all 16-bit monophonic recordings in .au format, sampled at 22kHz. We opted to use this dataset to assess the generalizability of our proposed pipeline on a smaller scale and a different corpus.
@@ 2.2 Methodologies
To extract perceptual features from the datasets, we employed three different techniques: Symbolic Knowledge, Deep Neural Networks, and Signal Processing.
@@
By utilizing symbolic knowledge, we created a new set of interpretable features that are derived from music theory. To accomplish this, we constructed a knowledge base in order to extract music features from each track’s chords. We selected Omnizart [29] from among many chord recognition models, based on its reported performance. Our knowledge base consists of chords as instances, and classes and subclasses from the Functional Harmony Ontology (FHO) [30]. These classes and subclasses describe the harmonic function of each chord in a music track. Using SPARQL queries, we retrieved the instances of certain classes to create useful features, such as the frequency of specific harmonic patterns (e.g., dominant-dominant-tonic), in the form of n-grams.
We employed deep neural networks (DNNs) to efficiently capture perceptual features. A VGG-style model was trained on a dataset titled ’Mid-level perceptual musical features’  [17]. For input, we used Mel-Frequency Cepstral Coefficients (MFCCs) extracted from the first 30 seconds of each audio file yielded a matrix $K\in\mathbb{R}^{S\times P}$, where $S$ is the number of segments and $P$ is the count of MFCCs. Subsequently, the trained model was applied to infer perceptual features from our target datasets (MTG-Jamendo, GTZAN), facilitating a comprehensive analysis and extraction of perceptual features from the music audio data.
We employed signal processing techniques to extract acoustically and musically meaningful features. These features were thoughtfully curated by music experts with the specific criterion of ensuring their accessibility and comprehensibility to human perception. If $x$ is the input audio file, we define our process as a function $E$, where $E(x)=x_{s}$, $x_{s}\in\mathbb{R}^{q}$ and $q\in\mathbb{N}$ ($q$ being the number of the extracted features).
Fig. 2: Feature importance with two different methods
@@ 3 Experiments
When dealing with a set of perceptual features, there exist two strategies for achieving explainable predictions of audio tags. The first approach involves employing explainable-by-design machine learning models , like Decision Trees. The second approach revolves around the utilization of post hoc black-box explanation techniques capable of elucidating the functioning of any machine learning model [31]. In our research, we adopt a combination of both approaches. We incorporate explainable-by-design models due to their capacity to provide readily understandable outputs when fed with human-interpretable features as input. Additionally, we leverage post hoc black-box explanation methods, such as SHAP values, to enhance our interpretability efforts. We provide the GitHub repository for the implementation [^1] https://github.com/vaslyb/PerceptibleMusicTagging.
@@ 3.1 Classification
In our quest to select a classifier, we conducted a thorough evaluation of various machine-learning models using the MTG-Jamendo dataset. We conducted experiments with six classifiers, namely the Decision tree, K-Nearest Neighbors algorithm, Extra Trees Classifier, Naive Bayes classifier, Random Forest, and XGBoost. These classifiers were configured with their default hyper-parameter settings, and they were fed with the 62 scalar perceptual features as input. Our evaluation led us to choose XGBoost as the most effective classifier. XGBoost is an ensemble machine learning classifier based on decision trees that employs a gradient-boosting framework.
Our XGBoost model achieved a ROC-AUC score of 0.729, slightly surpassing the baseline of 0.725 while falling slightly below the performance of state-of-the-art intelligent systems (0.83) [32]. Similar results were obtained when applying XGBoost to the multi-class classification task using the GTZAN dataset, where we achieved an accuracy of 0.79, showing a slight deviation from the state-of-the-art performance (0.84) [33, 34].
While our findings may not outperform the state-of-the-art deep learning models, the presence of interpretable model outputs serves the valuable purpose of identifying potential biases within the model. Furthermore, we achieve commendable results using models characterized by a limited number of trainable parameters.
@@ 3.2 Explanations
In this study, we chose XGBoost as the predictive modeling technique due to its interpretability. It enables us to gain insights into the significance of input features in the overall predictions. This characteristic has rendered it a valuable tool for improving our understanding of the classification process.
Overall, we utilized four different methods to extract feature importance. Firstly, we used the model’s built-in feature importance. Secondly, we computed a set of feature importance explanations using the permutation method, by randomly shuffling each feature and computing the change in performance. Thirdly, we computed SHAP values, which estimate how each feature contributes to the prediction based on notions from game theory. Lastly, we conducted ablation studies, as recommended in the literature [35], with different groups of features. We categorized our features into six groups based on their origin - midlevel, signal, and harmonic - and conducted experiments with all possible combinations of the groups, using XGBoost as our classifier.
Fig. 3: Ablation studies. The above figure for the MTG Jamendo dataset and the below figure for the GTZAN dataset
Fig. 4: Interpreting label Disco of the GTZAN dataset
@@ 4 Discussion
We obtained similar results in our ablation study (Fig. 3), where we found that signal processing features were the most useful, while harmonic features were the least useful for both datasets. Considering that harmonic features are extracted from the chord progression, this observation could be attributed to the quality of the chord recognition tool. Additionally, for the GTZAN dataset, only 30 sec for each track were given, leading to a limited number of chords per track and resulting in not having enough structural information. Harmonic features had a substantial impact only on predicting the Blues labels, a genre where chord progressions are distinct and easy to detect throughout a track. In terms of mid-level features, they appeared to be significant in improving the ROC-AUC metric for the Jamendo dataset. For both datasets, the best results were achieved by the combination of all features.
Fig. 5: Interpreting label Trailer of the MTG Jamendo dataset
We show examples of interpreting feature importance computed with SHAP values and XGBoost for predicting mood labels in the MTG Jamendo dataset. Specifically, we analyze one of the best-performing labels (ROC-AUC > 0.8) as shown in Fig. 5. For the highly-predictable Trailer mood predictions rely on low values of Length and high values of Spectral-Flux. A feature that is used for voice detection. This is aligned with the fact that trailers are usually shorter than songs and have voice-overs.
@@ 5 Conclusions and future work
We have created a specialized pipeline for music audio tagging with an emphasis on delivering results that are interpretable. Our approach incorporates established feature extraction methods alongside the introduction of a fresh set of features derived from the realm of music harmony concepts. Our experimental findings underscore the pivotal role of feature engineering within our proposed pipeline, as it exerts an influence on both performance and interpretability. It is imperative that each feature’s meaning remains comprehensible to humans, emphasizing the significance of this step.

### 23 Feb 2024 v3

@@ 1 Introduction
In these cases, the reliability of the training data dictates the reliability of the model, and without interpretability, it is difficult to detect flaws or biases in either data or models [12]. For music especially, where real-world data tend to be diverse, contrary to available datasets [13], it is essential to be able to detect and fix biases to make robust systems that can be applied in real-world settings [14]. As an extreme example, consider a mood tagging system that was trained only on pop and rock music. We cannot expect such a system to be generalized for all music genres, and it would not be possible to predict how it would behave if it were fed for instance classical music. Furthermore, we cannot expect an end-user to know the distribution of the training dataset, and the nuances of training ML models, thus such a system would not be applicable in practice, regardless of how well it performed on a test set. Contrarily, if the mood tagging system itself were interpretable, it would be easier for a user to understand why the system assigns specific tags to samples, making it more trustworthy and more likely to be applied in a real-world setting.
@@
In recent years, efforts have emerged aimed at addressing the interpretability challenge within music audio tagging systems [15]. Perceptual musical features are used that are readily comprehensible to human users [16, 17, 18, 19, 20]. These features encompass characteristics imbued with both musical and acoustic significance, such as tonality, rhythmic stability, or loudness, and they exhibit a strong correlation with emotions [21, 22]. Due to this inherent correlation, these features have demonstrated their efficacy in Music Emotion Recognition (MER) tasks. These characteristics are rooted in musical concepts and theoretical knowledge, as noted in a survey [23], though quantitative studies on them remain relatively scarce. In our research, we build upon these concepts and execute a structured approach for achieving interpretable music audio tagging (Fig. 1), harnessing the power of perceptual features. Our specific contributions can be succinctly summarized as follows:
  • In our study, we integrate three distinct methods for extracting perceptual musical features: signal processing, deep neural networks, and symbolic knowledge. Particularly noteworthy is our introduction of a novel set of features directly tied to music harmony. We demonstrate that these information sources are mutually complementary, significantly enhancing classifier performance across two datasets.
  • We interpret model predictions, using four different methodologies, SHAP [24], XGBoost [25] feature importance, permutation-based feature importance, and an ablation study.
@@ 2.1 Datasets
We employed two datasets in our experiments, the MTG-Jamendo dataset [26] and the GTZAN dataset [27]. The MTG-Jamendo dataset serves as an open resource for music audio tagging. This dataset was constructed using music accessible through the Jamendo platform, offered under Creative Commons licenses, and enriched with tags provided by content uploaders. Boasting a substantial scale, it encompasses more than 18,486 complete audio tracks, that we used, annotated with 56 tags that encompass various mood and theme categories. Our choice of this specific dataset was predicated on the anticipation that interpretability would offer heightened value in a task involving partially subjective aspects, such as mood detection [28].
The GTZAN dataset [27] is a collection comprising 1000 audio tracks of 30 seconds and features 10 genres, with 100 tracks representing each. The audio files are all 16-bit monophonic recordings in .au format, sampled at 22kHz. We opted to use this dataset to assess the generalizability of our proposed pipeline on a smaller scale and a different corpus.
@@ 2.2 Methodologies
To extract perceptual features from the datasets, we employed three different domains: Symbolic Knowledge, Deep Neural Networks, and Signal Processing.
@@
By utilizing symbolic knowledge, we created a new set of interpretable features that derive from music theory. To accomplish this, we constructed a knowledge base to extract music features from each track’s chords. We selected Omnizart [29] from among many chord recognition models, based on its reported performance. Our knowledge base consists of chords as instances, and classes and subclasses from the Functional Harmony Ontology (FHO) [30]. These classes and subclasses describe the harmonic function of each chord in a music track. Using SPARQL queries, we retrieved the instances of certain classes to create useful features, such as the frequency of specific harmonic patterns (e.g., dominant-dominant-tonic), in the form of n-grams. In total, we created a 32-dimensional feature vector $x_{h}\in{\mathbb{R}^{32}}$. First, we created Dominants and Subdominants by calculating the ratio between the sum of dominant and subdominant retrieved chords and the total number of chords, respectively. Additionally, by defining the function of a chord in any chord progression, we developed 30 more features based on bigrams and trigrams (e.g., sub-dom, sub-sub-dom). These features were defined by calculating the ratio of a specific bigram or trigram to the total number of bigrams or trigrams in the analysis, respectively. To construct these features, we also designated some chords as tonics - chords that follow a dominant without serving another function, and globs - major dominant chords that produce a great deal of harmonic tension.
We leveraged the power of deep neural networks (DNNs) to efficiently capture perceptual features from a dataset curated by music experts [17]. To build our DNN model, we used a vgg-ish architecture trained with multiple linear regression. We first extracted MFCCs with 40 coefficients, utilizing a window and FFT size of 2048, from the initial 15 seconds of each audio file. This procedure resulted in an input MFCC $K\in\mathbb{R}^{S\times P}$, where $S$ is the number of segments and $P$ is the number of MFCCs. The model’s output is a feature vector $x_{m}\in\mathbb{R}^{p}$, where $p\in\mathbb{N}$ corresponds to the number of extracted features. We trained our model on the "Mid Level Perceptual Music Features" dataset [17], comprising 5000 songs, each lasting 15 seconds in audio format. Each song is annotated by music experts with seven scalar perceptual features: Melodiousness, Articulation, Rhythmic Stability, Rhythmic Complexity, Dissonance, Tonal Stability, and Minorness.
Utilizing signal processing we extracted acoustically meaningful features. In order to choose features appropriate for our task a filtering process is essential having as a criterion their interpretability. If $x$ is the input audio file, we define our process as a function $E$, where $E(x)=x_{s}$, $x_{s}\in\mathbb{R}^{q}$ and $q\in\mathbb{N}$ ($q$ the number of the extracted features). For this approach, we used features extracted with a signal processing tool called Essentia [31]. The features were filtered out from three music experts, retaining only those that hold acoustical and musical significance and are easily understandable by humans. They ended up choosing 23 scalar features [^1] https://essentia.upf.edu/streaming_extractor_music.html#music-descriptors: Danceability, Loudness, Chords Changes Rate, Dynamic Complexity, Zero Crossing Rate, Chords Number Rate, Pitch Salience, Spectral Centroid, Spectral Complexity, Spectral Decrease, Spectral Energyband High, Spectral Energyband-Low, Spectral Energyband Middle High, Spectral Energyband Middle Low, Spectral Entropy, Spectral Flux, Spectral Rolloff, Spectral Spread, Onset Rate, Length, BPM, Beats Loud and Vocal-Instrumental.
The description of the features and the code for the implementation of the experiments can be found in the GitHub repository 222 https://github.com/vaslyb/PerceptibleMusicTagging . By concatenating $x_{h},x_{m},x_{s}$, and normalizing the values with a standard scaler, we end up with the feature vector used to train the models.
Fig. 2: Overall and permutation-based feature importance of the model trained on the MTG Jamendo dataset
@@ 3 Experiments
When dealing with a set of perceptual features, there exist two strategies for achieving explainable predictions of audio tags. The first approach involves employing explainable-by-design ML models , like Decision Trees. The second approach revolves around the utilization of post hoc black-box explanation techniques capable of elucidating the functioning of any ML model [32]. In our research, we adopt a combination of both approaches. We incorporate explainable-by-design models due to their capacity to provide readily understandable outputs when fed with human-interpretable features as input. Additionally, we leverage post hoc black-box explanation methods, such as SHAP values, to enhance our interpretability efforts.
@@ 3.1 Classification
After experimenting with a wide array of traditional ML classifiers, we ended up choosing XGBoost as the most effective classifier. XGBoost is an ensemble ML classifier based on decision trees that employs a gradient-boosting framework. Our model achieved a ROC-AUC score of 0.729, in the multi-label mood recognition task of the MTG Jamendo dataset, slightly surpassing the baseline of 0.725 [33] while falling slightly below the performance of state-of-the-art intelligent systems 0.769 [34]. Similar results were obtained when applying XGBoost to the multi-class classification task using the GTZAN dataset, where we achieved an accuracy of 0.79, showing a slight deviation from the state-of-the-art performance 0.84 [35, 36]. While our findings may not outperform the state-of-the-art deep learning models, the presence of interpretable model outputs serves the valuable purpose of identifying potential biases within the model. Furthermore, we achieve commendable results using models characterized by a limited number of trainable parameters.
@@ 3.2 Explanations
Interpretability played a pivotal role in the selection of XGBoost as the predictive modeling technique. It enabled us to gain insights into the significance of input features in the overall predictions. This characteristic has rendered it a valuable tool for improving our understanding of the classification process. Our study utilized four methods to extract feature importance from our XGBoost model, including the built-in feature importance, permutation method, SHAP values, and ablation studies.
Initially, we utilized the XGBoost feature importance, which involved assessing the weight of features based on their frequency in trees. By aggregating this information across all tags, we obtained a holistic overview of the features crucial to the task. Additionally, examining the weight of each specific tag allowed us to discern the importance of features on a tag-by-tag basis.
To further enhance our insights, we employed the permutation method to compute a second set of feature importance explanations. This involved randomly shuffling the values of each feature and observing the resulting degradation of the model’s score. This approach provided an overall perspective on the importance of features across all mood tags.
Subsequently, we incorporated SHAP values to extract feature importance, leveraging game theory concepts to estimate the contribution of each feature to predictions. Unlike the weight and permutation methods, SHAP values indicated whether features impacted the predictions positively, or negatively. The results were visualized through plots, showcasing the distribution of impacts each feature had on the model output, sorted by the sum of SHAP value magnitudes over all samples.
In line with recommended practices from the literature, we conducted an ablation study with different feature groups, categorized as Mid-level, Harmonic, and Signal based on their origin. By experimenting with all possible combinations of these groups, we gained valuable insights into the distinct impacts each set of features had on the model’s performance.
(a) MTG Jamendo
(b) GTZAN
Fig. 3: Ablation studies with different three groups of features: harmonic, signal, and midlevel on both datasets
Fig. 4: Interpretations of the prediction for the label Disco in the GTZAN dataset based on SHAP values and model’s feature importance.
@@ 4 Discussion
We obtained similar results in our ablation study (Fig. 3), where we found that signal processing features were the most useful, while harmonic features were the least useful for both datasets. Considering that harmonic features are extracted from the chord progression, this observation could be attributed to the quality of the chord recognition tool. Harmonic features had a substantial impact only on predicting the Blues labels, a genre where chord progressions are distinct and easy to detect throughout a track. In terms of mid-level features, they appeared to be significant in improving the ROC-AUC metric for the Jamendo dataset. For both datasets, the best results were achieved by the combination of all features.
Fig. 5: Interpretations of the prediction for the label Trailer in the MTG Jamendo dataset based on SHAP values and model’s feature importance
We show examples of interpreting feature importance computed with SHAP values and XGBoost for predicting mood labels in the MTG Jamendo dataset. Specifically, we analyze one of the best-performing labels (ROC-AUC > 0.8) as shown in Fig. 5. For the highly-predictable Trailer mood predictions rely on low values of Length and high values of Spectral-Flux, a feature that is used for voice detection. This is aligned with the fact that trailers are usually shorter than songs and have voice-overs.
@@
The above examples are meant to illustrate the value of interpretability for the audio tagging task. We can gain insights on a case-by-case basis which would not be possible in an end-to-end setup. Whether or not these insights are worth the additional work needed for feature extraction, and the slight deterioration in performance, depends on each application’s priorities. We argue that interpretability should be a priority for music audio tagging systems, as they suffer from label ambiguity [37], in addition to being at risk of the typical biases of MIR datasets [38].
@@ 5 Conclusions and future work
We have created a specialized pipeline for music audio tagging with an emphasis on delivering interpretable results. Our approach incorporates established feature extraction methods alongside the introduction of a fresh set of features derived from the realm of music harmony concepts. Our experimental findings underscore the pivotal role of feature engineering within our proposed pipeline, as it exerts an influence on both performance and interpretability. It is imperative that each feature’s meaning remains comprehensible to humans, emphasizing the significance of this step.
@@
Acknowledgments
We express our sincere gratitude to Christos Garoufis for his useful insights and thoughtful comments on our work. The research project is implemented in the framework of H.F.R.I call “Basic research Financing (Horizontal support of all Sciences)” under the National Recovery and Resilience Plan “Greece 2.0” funded by the European Union –NextGenerationEU(H.F.R.I. Project Number: $1$ Emotional Artificial Intelligence in Music Expression).